mathslover (mathslover):

Logarithm Question. Solve for x : \(\log_{\cfrac{3}{4}} \log_8 (x^2+7) + \log_{\cfrac{1}{2}} \log_{\cfrac{1}{4}} (x^2+7)^{-1} = - 2\)

mathslover (mathslover):

THis is what I have done yet : \(\boxed{\log_\cfrac{3}{4} \left[ \cfrac{1}{3}\log_2 (x^2 +7) \right] - \log_2 \left[ \cfrac{1}{2} \log_2 (x^2 + 7) \right] \\}\)

mathslover (mathslover):

The rule I used yet : \(\large \boxed {\log_{a^q} (x) = \cfrac{1}{q} \log_a x}\)

mathslover (mathslover):

@myininaya @mathmale

mathslover (mathslover):

Changing the base will work?

mathslover (mathslover):

Well, the second term can be simplified, I think : \(\log_\cfrac{1}{2} \cfrac{1}{2} + \log_\cfrac{1}{2}\left( \log_2 (x^2 + 7) \right) \)

Parth (parthkohli):

Let \(\log_2(x^2 + 7) = t\)

mathslover (mathslover):

Okay, 1 + \(\log_\cfrac{1}{2} t \)

Parth (parthkohli):

No!

mathslover (mathslover):

So, I get : \(\log_\cfrac{3}{4} \left(\cfrac{1}{3} t \right) + 1 + \log_\cfrac{1}{2} t = -2 \)

Parth (parthkohli):

The second term is log base 2 right?

Parth (parthkohli):

NVM.\[\boxed{\log_\cfrac{3}{4} \left[ \cfrac{1}{3}\log_2 (x^2 +7) \right] - \log_2 \left[ \cfrac{1}{2} \log_2 (x^2 + 7) \right] \\}\]Simplify the second term to \(-(1 + \log_2 t) \)

Parth (parthkohli):

And the first term to\[\log_{3/4} 1/3 + \log_{3/4} t\]

Parth (parthkohli):

Yeah, whoops.

Parth (parthkohli):

The second term is \(-\left(-1 + \log_2 t\right)\)

mathslover (mathslover):

Sorry, my net got down... Well yeah, I happened to make some confusion here.

Parth (parthkohli):

\[\log_{3/4} (1/3) + \log_{3/4}(t) + 1 - \log_2 t = 2\]Yeah, we're almost done.

Parth (parthkohli):

Write \(1/3\) as \(3^{-1}\).

Parth (parthkohli):

@mathslover Can you continue?

Parth (parthkohli):

Too many things to follow. Wait, I'll try to do this on paper and will return to you.

Parth (parthkohli):

\[\log_{3/4}(t/3) - \log_2 (t/2)=-2\]\[\log_{3/4}t - \log_{3/4}3 - \log_2 t + 1 = -2\]\[\log_{3/4} t - \log_2 t = -3 + \log_{3/4} 3\]

Parth (parthkohli):

I would try to get a common base on the LHS, but... lol

OpenStudy (anonymous):

Yeah make it into a single fraction, and it's simple from there, kind of...you'll get a complex number, I'm sure.

mathslover (mathslover):

\(-\log_\cfrac{3}{4} 3 + \log_2 t . \log_\cfrac{3}{4} 2 - \log_2 t = -3 \)

Parth (parthkohli):

Use `\cfrac` only for continued fractions...

mathslover (mathslover):

(guys, my net is not working properly, so, sorry for late responses) And fromt he above equation : \(\color{blue}{\text{Originally Posted by}}\) @mathslover \(-\log_\cfrac{3}{4} 3 + \log_2 t . \log_\cfrac{3}{4} 2 - \log_2 t = -3 \) \(\color{blue}{\text{End of Quote}}\) We can continue like : \(-\log_\cfrac{3}{4} 3 + \log_2 t ( \log_\cfrac{3}{4} 2 - 1 ) = -3\) or \(\log_2 t (\log_\cfrac{3}{4} 2 - 1) = \log_\cfrac{3}{4} 3 -3\)

Parth (parthkohli):

Ah, nice :)

mathslover (mathslover):

\(\log_2 t \left( \log_\cfrac{3}{4} \cfrac{8}{3}\right) = \log_\cfrac{3}{4} \left(\cfrac{3}{4}\right)^{-3} + \log_\cfrac{3}{4} 3 \)

mathslover (mathslover):

Seems like it is almost done..

mathslover (mathslover):

\(\log_2 t \times \log _ {3/4} \cfrac{8}{3} = \log_{3/4} \cfrac{3^{-2}}{4^{-3}} \\ \log_2 t \times \log_{3/4} \cfrac{8}{3} = \log_{3/4} \left(\cfrac{3^{-1}}{2^{-3}}\right)^2 \\ \log_2 t \times \log_{3/4} \cfrac{8}{3} =2 \log_{3/4} \cfrac{8}{3} \)

mathslover (mathslover):

\(\log_2 t \times \cancel{\log_{3/4} \cfrac{8}{3}} = 2\cancel{\log_{3/4}{ \cfrac{8}{3}}} \) \(\log_2 t = 2\)

Parth (parthkohli):

Phew, at least I got close to solving it :)

mathslover (mathslover):

Wooh... t = 4 putting t = \(\log_2 {x^2 + 7}\) \(\log_2 {x^2 + 7} = 4\) \(x^2 + 7 = 16\) \(x^2 = 9\) Thus \(x = \pm 3\)

mathslover (mathslover):

Yeah @ParthKohli ... Thanks for your help.